# Refuting Perfect Matchings in Spectral Expanders is Hard

### v1

@@ 1 Introduction
Perhaps the most fundamental problem in computation is to provide an answer to the question asked by Cook and Reckchow in their seminal paper [cook1979relative] – “Given a true statement $A$, is there a short proof of the claim that $A$ is true?” In trying to answer this question, we must first describe what constitutes a valid proof. That is, we must describe the language in which the proof is written (axioms), and the rules for checking it (the verifier). Each set of rules for writing and checking a proof defines a proof system. Therefore, a precise restatement of the question above is the following: “Given a true statement $A$ and a proof system $S$, what is the length of a shortest $\pi\in S$ that proves $A$?” If we could show that there exists a proof system $S$, such that for any true statement $A$, the length of the shortest proof in $S$ is upper bounded by some polynomial in the length of $A$, it would imply that $\mathsf{coNP}=\mathsf{NP}$ and consequently $\mathsf{P}=\mathsf{NP}$. Conversely, if we could show large proof size lower bounds for some $A$ in all proof systems, it would lead to a formal proof of the widely believed conjecture that $\mathsf{P}\neq\mathsf{NP}$. Unfortunately, such lower bounds for arbitrary proof systems are out of reach. As an intermediate step, the research community has invested a significant effort in proving lower bounds for increasingly expressive proof systems (e.g. see [abascal2021strongly, alekhnovich2001lower, atserias2020Size, buss1999linear, conneryd2023graph, de2023clique, impagliazzo1999lower, kothari2017SumOfSquares, potechin:LIPIcs.CCC.2020.38, raz2008elusive, razborov1998lower, schoenebeck2008linear]).
@@
In this work, we focus on the algebraic and semi-algebraic proof systems of polynomial calculus (PC) and sum of squares (SoS). In algebraic proof systems, we are given a set $\mathcal{Q}=\{q_{i}(\mathbf{x})\text{ }|\text{ }i\in[m]\}$ of $m$ polynomial equations [^1] Semi-algebraic proof systems also allow for inequalities but we will not deal with inequality constraints in this paper. over $n$ variables $\mathbf{x}=\{x_{1},\dots,x_{n}\}$. In PC, the coefficientcs can be over an arbitrary field $\mathbb{F}$ and in the SoS the coefficients are over the reals. We say a proof $\pi$ is a refutation of $\mathcal{Q}$, if it is a proof of the claim (in the specified language) that there exists no assignment of $\mathbf{x}\in\mathbb{F}^{n}$ that satisfies all the polynomial equations in $\mathcal{Q}$. In PC and SoS, the proof $\pi$ is itself expressed as a sequence of polynomials. Two common measures of the complexity of the proof are the length of the sequence and the largest degree of the proof polynomials that refute $Q$ (see Section 2.1 for more formal details). Trade-offs between the two are well known; in particular, a lower bound on the degree implies a comparatively strong lower bound on the size of the proof. Therefore, in this work we focus on the former. We denote the smallest maximum degree over all proofs that refute $\mathcal{Q}$ in PC and SoS, with $\text{Deg}\left(Q\vdash_{\texttt{PC}_{F}}\bot\right)$ and $\text{Deg}\left(Q\vdash_{\texttt{SOS}}\bot\right)$, respectively. One motivation for proving lower bounds for algebraic proof systems, as opposed to propositional proof systems, is that often they imply lower bounds for a broad family of related algorithms for solving combinatorial optimisation problems. Similarly, upper bounding the proof length has led to the fruitful discovery of many efficient algorithms. The SoS proof system is of particular interest because of its close connection to the sum-of-squares hierarchy of semi-definite programming. We refer the reader to the survey by Fleming, Kothari, and Pitassi [fleming2019semialgebraic] for more details about the connections between the semi-algebraic proof systems and combinatorial optimisation.
@@
In this work, we study the complexity of refuting perfect matchings in PC and SoS. Apart from being a natural problem in its own right, perfect matchings are also related to the pigeon hole principle [maciel2000new, pitassi1993exponential, raz2004resolution, razbarov2002pgp, razborov2003resolution] and Tseitin formula [filmus2013towards, galesi2019bounded, glinskih2017satisfiable, grigoriev2001linear], two well studied formulae in proof complexity. Assuming at most one pigeon fits in a single hole, the pigeon hole principle says $m$ pigeons cannot fit in $n<m$ holes. If we construct the complete bipartite graph with the left vertices as $m$ pigeons and the right vertices as $n<m$ holes, proving the pigeon hole principle amounts to proving that such bipartite graph does not have a perfect matching. There are other formulations of the pigeon hole principle (see the survey by Razborov [razbarov2002pgp]), and almost all of them have short proofs in the sum of squares proof system. In contrast, Tseitin formulae is known to require long proofs. The Tseitin formula over a graph claims that there is a spanning subgraph in which every vertex has odd degree. If a graph has a perfect matching, then the subgraph described by the matching ensures that every vertex has odd degree. However, formally refuting Tseitin formulae for expander graphs with odd number of vertices, in the SoS proof system, requires degree linear in the number of vertices in the graph [grigoriev2001linear]. Given its close connections to the pigeon-hole and Tseitin, and the different behaviour of the two formulae, it is natural to determine the complexity of refuting perfect matchings for non-bipartite graphs.
@@
For every $e\in E$, the equation $x_{e}(1-x_{e})=0$ restricts the domain of the above variables to bits. In plain words, $\text{Card}(G,\mathbf{b})$ denotes the claim that there exists a spanning subgraph $G^{\prime}\subseteq G$ such that a vertex $v\in V(G)$ has $b_{v}$ edges incident to it in $G^{\prime}$. Note if there was an assignment of variables $(x_{e})_{e\in E}$ that satisfies all equations in $\text{Card}(G,\mathbf{1})$, where $\mathbf{1}=(1,\ldots,1)\in\mathbb{F}^{n}$, it would imply that the graph $G$ has a perfect matching (given by the edges corresponding to variables with assignment 1). Therefore, we define $\text{PM}(G):=\text{Card}(G,\mathbf{1})$. When $|V|$ is odd, $G$ trivially does not contain a perfect matching. How difficult is it to refute $\text{PM}(G)$ in this case? In recent work, Austrin and Risse [Austrin_2022] showed that refuting $\text{PM}(G)$, in the Polynomial Calculus and Sum-of-Squares system, in the case $G$ is a random $d$-regular graphs with an odd number of vertices typically requires proofs with degree $\Omega\left(n/\log n\right)$. They conjecture (see [Austrin_2022, Section 6]) that the hardness results should also apply to general expander graphs but leave showing so as an open problem. In this work, we verify this by extending their result to all $d$-regular spectral expanders, that is, $d$-regular graphs with a mild condition on the spectral gap. In fact, similar to Austrin and Risse, we reduce the hardness of refuting $\text{Card}(G,\mathbf{t})$, for any odd value $t$, to the hardness of refuting $\text{Card}(G,\mathbf{1})$. As another special case, this answers the even-colouring case when $t=d/2$ is odd, a problem posed by Buss and Nordström [buss2021proof, Open Problem 7.7], which asks, “Are even colouring formulas over expander graphs hard for polynomial calculus over fields of characteristic distinct from 2 ?” Formally, we prove the following (for the definition of $(n,d,\lambda)$-graphs see Section 2.2).
@@
We follow the overall approach of Austrin and Risse [Austrin_2022]. Very briefly, the strategy is to obtain an affine restriction (see Definition 2.6) $\text{Card}(G,\mathbf{t})|_{\rho}\equiv\text{PM}(H)$ where $H$ is some graph for which refuting $\text{PM}(H)$ requires large degree. An example of such $H$ is given by Buss, Grigoriev, and Impagliazzo [buss1999linear]. We now describe how to find such a restriction in more details: Using a result of Draganić, Krivelevich, and Nenadov [draganic22rolling], we show that $H$ topologically embeds into a given expander graph $G$ with $\lambda\leq\varepsilon d$ for some universal small constant $\varepsilon\in(0,1)$, such that all paths corresponding to the embedding have odd length. The main technical ingredient of Austrin and Risse is also similar embedding theorem, but one that critically relies on the underlying graph being random. As a random $d$-regular graph is with high probability an $(n,d,\lambda)$-graph with $\lambda=\Theta(\sqrt{d})$ (see [tikhomirov2016spectralgapdenserandom, Theorem A]), our embedding theorem subsumes that of [Austrin_2022]. Moreover, we show that one can find such an embedding so that the subgraph of $G$ induced by vertices which are not part of the embedding has a perfect matching. This allows us to use the restriction argument to transfer the hardness of $\text{PM}(H)$ into the hardness of $\text{PM}(G)$. To extend this to hardness of $\text{Card}(G,\mathbf{t})$ for an odd $3\leq t\leq d$, it suffices to show that the graph $G^{\prime}$ obtained from $G$ by removing all edges that participate in the embedding and the matching contains a $(t-1)$-regular spanning subgraph. Austrin and Risse achieve this using the contiguity property of random regular graphs. Instead, we provide a significantly simpler and shorter argument based on Tutte’s criterion.
The rest of the document as is structured as follows. In section 2, we describe the requisite background from graph theory and proof complexity required. In Section 3, we describe the machinery for finding a desired topological embedding. In Section 4, we prove conditions under which the residual graph has a perfect matching or, more generally, a $(t-1)$-regular spaning subgraph. In Section LABEL:sec:main-proof, we use the tools from the previous sections to prove Theorem 1.1. In Section LABEL:sec:related-work we briefly discuss a few other lower bounds using embeddings in proof complexity, and conclude with some future directions.
@@ 2.1 Proof Complexity Preliminaries
Given a set of $m$ polynomial equality constraints $\mathcal{Q}$ over the reals, a Sum of Squares (SoS) refutation is a set of polynomials $\pi=(t_{1},\dots,t_{m};s_{1},\dots,s_{a})$ such that
@@
Note that $s_{i}^{2}(x)\geq 0$ for any $x$ by definition. Therefore, if there were to exist some $x^{\star}$ such that $p_{i}(x^{\star})=0$ for all $p_{i}\in\mathcal{Q}$, then $\sum_{i\in[m]}t_{i}(x^{\star})p_{i}(x^{\star})=0$. This would imply that $h\geq 0$, but our proof $\pi$ shows that $h=-1$. Therefore, the existence of the set of polynomials $\pi$ act as a formal proof of the claim that the set of polynomials equations in $\mathcal{Q}$ is unsatisfiable.
@@
Polynomial Calculus (PC) is a dynamic version of the static Nullstellensatz proof system [fleming2019semialgebraic, Section 1.3] based on the following inference rules.
@@
A Polynomial Calculus (PC) refutation of $\mathcal{Q}$ over is a set of polynomials $\pi=t_{1},\dots,t_{l}$ such that $t_{l}=1$, and for each $i\neq l$, either (1) $t_{i}\in\mathcal{Q}$, or (2) $t_{i}$ is derived from $\{t_{j}\}_{j<i}$ using the above rules. The degree of the proof is given by $\text{Deg}\left(\pi\right)=\max_{i\in l}\text{Deg}\left(t_{i}\right)$. If we let $\Pi$ denote the set of all PC refutations of $\mathcal{Q}$, then
@@
To ensure boolean variables, we assume the axioms $\mathcal{Q}$ always contain the equations $x_{i}^{2}-x_{i}=0$ and $\bar{x}^{2}_{i}-\bar{x}_{i}=0$ for all $i\in[n]$. Equivalently, we can also just work in the ring $\mathbb{F}[x_{1},\dots,\bar{x}_{n}]/(x_{1}^{2}-x_{1},\dots,\bar{x}_{n}^{2}-\bar{x}_{n})$ of multi-linear polynomials. Multi-linearity implies that the degree of any proof can be at most $n$ i.e a proof of degree $\Omega\left(n\right)$ is the largest lower bound one can hope to achieve. Additionally, we will also assume that $01-x_{i}-\bar{x}_{i}=$ 0 is also included in $\mathcal{Q}$, for all $i\in[n]$, which ensures that the bar elements are bit complements of the non-bar elements. The following lemma is by Buss, Grigoriev, and Impagliazzo [buss1999linear] and gives an instance where perfect matching is hard to refute in the worst case.

### v2

@@ 1 Introduction
Perhaps the most fundamental problem in computation is to provide an answer to the question asked by Cook and Reckchow in their seminal paper [cook1979relative] – “Given a true statement $A$, is there a short proof of the claim that $A$ is true?” In trying to answer this question, we must first describe what constitutes a valid proof. That is, we must describe the language in which the proof is written (axioms), and the rules for checking it (the verifier). Each set of rules for writing and checking a proof defines a proof system. Therefore, a precise restatement of the question above is the following: “Given a true statement $A$ and a proof system $S$, what is the length of a shortest $\pi\in S$ that proves AA?” If we could show that there exists a proof system $S$, such that for any true statement $A$, the length of the shortest proof in $S$ is upper bounded by some polynomial in the length of $A$, it would imply that $\mathsf{coNP}=\mathsf{NP}$ and consequently the polynomial hierarchy collapes to $\mathsf{NP}$. Conversely, if we could show large proof size lower bounds for some $A$ in all proof systems, it would lead to a formal proof of the widely believed conjecture that $\mathsf{P}\neq\mathsf{NP}$. Unfortunately, such lower bounds for arbitrary proof systems are out of reach. As an intermediate step, the research community has invested a significant effort in proving lower bounds for increasingly expressive proof systems (e.g. see [abascal2021strongly, alekhnovich2001lower, atserias2020Size, buss1999linear, conneryd2023graph, de2023clique, impagliazzo1999lower, kothari2017SumOfSquares, potechin:LIPIcs.CCC.2020.38, raz2008elusive, razborov1998lower, schoenebeck2008linear]).
@@
In this work, we focus on the algebraic and semi-algebraic proof systems [^1] Similar to the work of Austrin and Risse [Austrin_2022], our lower bounds also immediately extend to bounded depth Frege proof systems. However, the key technical contribution in this work is the graph theoretic techniques proposed for embedding a hard instance into the host graph. Provided we are able to find an appropriate hard instance (which is known to exist for Frege proof systems), our techniques are independent of the underlying proof systems. Thus, we restrict our preliminaries to PC and SoS to keep our preliminaries stand-alone, and yet concise. of polynomial calculus (PC) and sum of squares (SoS). In algebraic proof systems, we are given a set $\mathcal{Q}=\{q_{i}(\mathbf{x})\text{ }|\text{ }i\in[m]\}$ of $m$ polynomial equations [^2] Semi-algebraic proof systems also allow for inequalities but we will not deal with inequality constraints in this paper. over $n$ variables $\mathbf{x}=\{x_{1},\dots,x_{n}\}$. In PC, the equations can be over an arbitrary, but fixed field $\mathbb{F}$, and in the SoS the coefficients are over the reals. We say a proof $\pi$ is a refutation of $\mathcal{Q}$, if it is a proof of the claim (in the specified language) that there exists no assignment of $\mathbf{x}\in\mathbb{F}^{n}$ that satisfies all the polynomial equations in $\mathcal{Q}$. In PC and SoS, the proof $\pi$ is itself expressed as a sequence of polynomials. Two common measures of the complexity of a semi-algebraic proof are size (the number of monomials appearing in the proof) and the degree (the largest degree of the proof polynomials that refute $Q$). Trade-offs between the two are known. If $\mathcal{Q}$ is the indexed set of polynomials with degree at most $k$, and if $d$ is the degree of the SoS refutation; then Atserias and Hakoniemi [atserias2020Size, Corollary 1] show that the size $s$ of the refutation is related as $s\geq\exp\left(\frac{(d-k-4)^{2}}{32(n+1)}\right)$. In this work, to prove hardness of refutations in the SoS proof system we show degree lower bounds of $d\in\Omega(n/\log n)$. By the above relation, this implies exponential size lower bounds in $n$ as well. Similar size-degree trade-offs are also known for PC [impagliazzo1999lower], and thus it suffices for us to just show degree lower bounds to get exponential size lower bounds. We denote the smallest maximum degree over all proofs that refute $\mathcal{Q}$ in PC and SoS, with $\text{Deg}\left(Q\vdash_{\texttt{PC}_{F}}\bot\right)$ and $\text{Deg}\left(Q\vdash_{\texttt{SOS}}\bot\right)$, respectively. One motivation for proving lower bounds for algebraic proof systems, as opposed to propositional proof systems, is that often they imply lower bounds for a broad family of related algorithms for solving combinatorial optimisation problems. Similarly, upper bounding the proof length has led to the fruitful discovery of many efficient algorithms. The SoS proof system is of particular interest because of its close connection to the sum-of-squares hierarchy of semi-definite programming. We refer the reader to the survey by Fleming, Kothari, and Pitassi [fleming2019semialgebraic] for more details about the connections between the semi-algebraic proof systems and combinatorial optimisation.
@@
In this work, we study the complexity of refuting perfect matchings in PC and SoS. Apart from being a natural problem in its own right, perfect matchings are also related to the pigeon hole principle [maciel2000new, pitassi1993exponential, raz2004resolution, razbarov2002pgp, razborov2003resolution] and Tseitin formula [filmus2013towards, galesi2019bounded, glinskih2017satisfiable, grigoriev2001linear], two well studied formulae in proof complexity. Assuming at most one pigeon fits in a single hole, the pigeon hole principle says $m$ pigeons cannot fit in $n<m$ holes. If we construct the complete bipartite graph with the left vertices as $m$ pigeons and the right vertices as $n<m$ holes, proving the pigeon hole principle amounts to proving that such bipartite graph does not have a perfect matching. There are other formulations of the pigeon hole principle (see the survey by Razborov [razbarov2002pgp]), and almost all of them have short proofs in the sum of squares proof system. In contrast, Tseitin formulae are known to require long proofs. The Tseitin formula over a graph claims that there is a spanning subgraph in which every vertex has odd degree. If a graph has a perfect matching, then the subgraph described by the matching ensures that every vertex has odd degree. However, formally refuting Tseitin formulae for expander graphs with an odd number of vertices, in the SoS proof system, requires degree linear in the number of vertices in the graph [grigoriev2001linear]. Given its close connections to the pigeon-hole and Tseitin, and the different behaviour of the two formulae, it is natural to determine the complexity of refuting perfect matchings for non-bipartite graphs.
@@
For every $e\in E$, the equation $x_{e}(1-x_{e})=0$ restricts the domain of the above variables to bits. In plain words, $\text{Card}(G,\mathbf{b})$ denotes the claim that there exists a spanning subgraph $G^{\prime}\subseteq G$ such that a vertex $v\in V(G)$ has $b_{v}$ edges incident to it in $G^{\prime}$. Note if there was an assignment of variables $(x_{e})_{e\in E}$ that satisfies all equations in $\text{Card}(G,\mathbf{1})$, where $\mathbf{1}=(1,\ldots,1)\in\mathbb{F}^{n}$, it would imply that the graph $G$ has a perfect matching (given by the edges corresponding to variables with assignment 1). Therefore, we define $\text{PM}(G):=\text{Card}(G,\mathbf{1})$. When $|V|$ is odd, $G$ trivially does not contain a perfect matching. How difficult is it to refute $\text{PM}(G)$ in this case? In recent work, Austrin and Risse [Austrin_2022] showed that refuting $\text{PM}(G)$, in the Polynomial Calculus and Sum-of-Squares system, in the case $G$ is a random $d$-regular graphs with an odd number of vertices typically requires proofs with degree $\Omega\left(n/\log n\right)$. They conjecture (see [Austrin_2022, Section 6]) that the hardness results should also apply to general expander graphs but leave showing so as an open problem. In this work, we verify this by extending their result to all $d$-regular spectral expanders, that is, $d$-regular graphs with a mild condition on the spectral gap. In fact, similar to Austrin and Risse, we reduce the hardness of refuting $\text{Card}(G,\mathbf{t})$, where $\mathbf{t}=(t,\ldots,t)$, for any odd value $t$, to the hardness of refuting $\text{Card}(G,\mathbf{1})$, where $\mathbf{1}=(1,\ldots,1)$. As another special case, this answers the even-colouring case when $t=d/2$ is odd, a problem posed by Buss and Nordström [buss2021proof, Open Problem 7.7], which asks, “Are even colouring formulas over expander graphs hard for polynomial calculus over fields of characteristic distinct from 2 ?” Formally, we prove the following (for the definition of $(n,d,\lambda)$-graphs see Section 2.2).
@@
We follow the overall approach of Austrin and Risse [Austrin_2022]. Very briefly, the strategy is to obtain an affine restriction (see Definition 2.6) $\text{Card}(G,\mathbf{t})|_{\rho}\equiv\text{PM}(H)$ where $H$ is some graph for which refuting $\text{PM}(H)$ requires large degree. An example of such $H$ is given by Buss, Grigoriev, and Impagliazzo [buss1999linear]. We now describe how to find such a restriction in more details: Using a result of Draganić, Krivelevich, and Nenadov [draganic22rolling], we show that $H$ topologically embeds into a given expander graph $G$ with $\lambda\leq\varepsilon d$ for some universal small constant $\varepsilon\in(0,1)$, such that all paths corresponding to the embedding have odd length. The main technical ingredient of Austrin and Risse is also a similar embedding theorem, albeit a significantly more complicated one. Moreover, we show that one can find such an embedding so that the subgraph of $G$ induced by vertices which are not part of the embedding has a perfect matching. This allows us to use the restriction argument to transfer the hardness of $\text{PM}(H)$ into the hardness of $\text{PM}(G)$. To extend this to hardness of $\text{Card}(G,\mathbf{t})$ for an odd $3\leq t\leq d$, it suffices to show that the graph $G^{\prime}$ obtained from $G$ by removing all edges that participate in the embedding and the matching contains a (t1)(t-1)-regular spanning subgraph. Austrin and Risse achieve this using the contiguity property of random regular graphs (and hence their hardness result for $\text{Card}(G,\mathbf{t})$ critically relies on randomness). Instead, we provide a significantly simpler and shorter argument based on Tutte’s criterion. As a random $d$-regular graph is with high probability an $(n,d,\lambda)$-graph with $\lambda=\Theta(\sqrt{d})$ (see [tikhomirov2016spectralgapdenserandom, Theorem A]), our embedding theorem readily applies in the context of [Austrin_2022].
The rest of the document is structured as follows. In section 2, we describe the requisite background from graph theory and proof complexity. In Section 3, we describe the machinery for finding a desired topological embedding. In Section 4, we prove conditions under which the residual graph has a perfect matching or, more generally, a (t1)(t-1)-regular spaning subgraph. In Section 5, we use the tools from the previous sections to prove Theorem 1.1. In Section 6 we briefly discuss a few other lower bounds using embeddings in proof complexity, and conclude with some future directions.
@@ 2.1 Proof Complexity Preliminaries
Given a set of $m$ polynomial equality constraints $\mathcal{Q}$ over the reals, a Sum of Squares (SoS) refutation is a sequence of polynomials $\pi=(t_{1},\dots,t_{m};s_{1},\dots,s_{a})$ such that
@@
Note that $s_{i}^{2}(x)\geq 0$ for any $x$ by definition. Therefore, if there were to exist some $x^{\star}$ such that $p_{i}(x^{\star})=0$ for all $p_{i}\in\mathcal{Q}$, then $\sum_{i\in[m]}t_{i}(x^{\star})p_{i}(x^{\star})=0$. This would imply that $h\geq 0$, but our proof $\pi$ shows that h=1h=-1. Therefore, the existence of the set of polynomials $\pi$ act as a formal proof of the claim that the set of polynomial equations in $\mathcal{Q}$ is unsatisfiable.
@@
Polynomial Calculus (PC) is a dynamic version of the static Nullstellensatz proof system [fleming2019semialgebraic, Section 1.3] operating over an arbitrary but fixed field, based on the following inference rules.
@@
A Polynomial Calculus (PC) refutation of $\mathcal{Q}$ over $\mathbb{F}$ is a sequence of polynomials $\pi=(t_{1},\dots,t_{l})$ such that $t_{l}=1$, and for each $i\neq l$, either (1) $t_{i}\in\mathcal{Q}$, or (2) $t_{i}$ is derived from $\{t_{j}\}_{j<i}$ using the above rules. The degree of the proof is given by $\text{Deg}\left(\pi\right)=\max_{i\in l}\text{Deg}\left(t_{i}\right)$. If we let $\Pi$ denote the set of all PC refutations of $\mathcal{Q}$, then
@@
To ensure Boolean variables, we assume the axioms $\mathcal{Q}$ always contain the equations $x_{i}^{2}-x_{i}=0$ and $\bar{x}^{2}_{i}-\bar{x}_{i}=0$ for all $i\in[n]$. Equivalently, we can also just work in the ring $\mathbb{F}[x_{1},\dots,\bar{x}_{n}]/(x_{1}^{2}-x_{1},\dots,\bar{x}_{n}^{2}-\bar{x}_{n})$ of multi-linear polynomials. Multi-linearity implies that the degree of any proof can be at most $n$ i.e a proof of degree $\Omega\left(n\right)$ is the largest lower bound one can hope to achieve. Additionally, we will also assume that $1-x_{i}-\bar{x}_{i}=0$ is also included in $\mathcal{Q}$, for all $i\in[n]$, which ensures that the bar elements are bit complements of the non-bar elements. The following lemma is by Buss, Grigoriev, and Impagliazzo [buss1999linear] and gives an instance where perfect matching is hard to refute in the worst case.